OpenStudy (anonymous):

Two vertices of a right isosceles triangle are (1,1) and (2,0). Which of the following could be the coordinates of the third vertex of this triangle?

OpenStudy (anonymous):



OpenStudy (anonymous):

This is like Theorem of pythagoras. Finding the slant 

OpenStudy (anonymous):

Finding the slant AC is the hypothenuse, by \[\frac{ 0-1 }{ 2-1 }= \frac{ -1 }{ 1 }=-1\]

OpenStudy (anonymous):

Given that it is an isosceles triangle, it means that sides AB and BC have the same length or slant. Let the coordinates of B be (x,y) in an equation we can make \[\frac{ 1-y }{ 1-x }=\frac{ 0-y }{ 2-x }\]

OpenStudy (anonymous):

Oops, guess I am wrong. Never mind with my work.